OpenStudy (anonymous):

Find the value of the Limit : \[\color{blue}{\lim_{x \rightarrow \infty}( \frac{x + \sin(x)}{x + \cos(x)})}\]

OpenStudy (anonymous):

just try to apply L'hospital Rule

OpenStudy (anonymous):

Okay: After applying L' Hospital Rule for first time, I am getting like this: \[\lim_{x \rightarrow \infty} (\frac{1 + \cos(x)}{1 - \sin(x)})\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} \frac{ 1+\cos x }{ 1-\sin x }\]

OpenStudy (anonymous):

is this is write?

OpenStudy (anonymous):

Yes it is...

OpenStudy (anonymous):

then rationalize it,i think you can get the answer

OpenStudy (anonymous):

After taking L'Hospital you are saying to apply Rationalization ??? @dirtymind

OpenStudy (anonymous):

I think its 1..we have to use expansions of sinx and cosx

OpenStudy (anonymous):

Or you are suggesting other method by using Rationalization???

OpenStudy (anonymous):

even i say rationlize it

OpenStudy (anonymous):

Can you explain properly @AbhimanyuPudi ..

OpenStudy (anonymous):

by multiplying and dividing by 1-sin x

OpenStudy (anonymous):

yes rationalize it by 1+sinx

OpenStudy (anonymous):

1 + sin(x) @uzumakhi

OpenStudy (anonymous):

yes apply it

OpenStudy (anonymous):

Let us try as you are saying ..

OpenStudy (anonymous):

Substitute these formulae: sinx = x - x^3/3! + x^5/51.... cosx = 1 - x^2/2! + x^4/4!...

OpenStudy (anonymous):

thanks @waterineyes

OpenStudy (anonymous):

why wont you use sandwich ?

OpenStudy (anonymous):

\[\frac{(1 + \cos(x)) \cdot (1 + \sin(x))}{\cos^2(x)} \implies \frac{1 + \sin(x) + \cos(x) + \sin(x) \cdot \cos(x)}{\cos^2(x)}\]

OpenStudy (anonymous):

I don't know the answer properly..

OpenStudy (anonymous):

you can bound it and use sandwich \[\frac{ x-1 }{ x+1 } < .. < \frac{ x+1 }{ x-1 }\]

OpenStudy (anonymous):

is there anything wrong here?

OpenStudy (anonymous):

What is Sandwich? Ca you explain it I am hearing this word in Limits for first time...

OpenStudy (anonymous):

now take sin^2x at denominator of each function and make 1/infinity=0

OpenStudy (anonymous):

first of all .. using this sandwich or squeeze gives 1 as the answer.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Squeeze_theorem

OpenStudy (anonymous):

How can I take that @dirtymind

hartnn (hartnn):

may i introduce a simpler approach ?

OpenStudy (anonymous):

if you have limf(x) and you can bound it by g(x) and h(x) g(x) <= f(x) <= h(x) if limg(x) = limh(x) = L then limf(x) = L

OpenStudy (anonymous):

Always @hartnn

hartnn (hartnn):

put x=1/y when x->infinity, y->0

OpenStudy (anonymous):

so as you see lim[ (x-1) / (x+1) ] = lim [ (x+1)/ (x-1) ] = 1 so your limit is 1 .. done

OpenStudy (anonymous):

I am getting your point @Coolsector ..

hartnn (hartnn):

i also liked the sandwitch approach by coolsector

OpenStudy (anonymous):

Yeah I got it @Coolsector

OpenStudy (anonymous):

now it looks like that i talked to myself:P

OpenStudy (anonymous):

Thanks for your precious help...

OpenStudy (anonymous):

yw

hartnn (hartnn):

u could also get it wothout sandwitch

OpenStudy (anonymous):

By using your idea @hartnn ??

hartnn (hartnn):

u know what is lim x->0 x sin(1/x) = ??

OpenStudy (anonymous):

0...

hartnn (hartnn):

0

OpenStudy (anonymous):

Yeah I said the same..

hartnn (hartnn):

sin(1/x) always lies between -1 and 1 and so as x approaches zero, 0<=|x sin(1/x)| <= |x| but since x->0, |x sin(1/x)| must approach zero also. similarly for cos

OpenStudy (anonymous):

I liked all the three discussed approaches to this problem...I wish i could give 3 medals to @hartnn , @Coolsector , @AbhimanyuPudi

OpenStudy (anonymous):

Then what?? I am not getting properly... @hartnn

hartnn (hartnn):

ok, u put y=1/x in original limit , what u get ?

OpenStudy (anonymous):

Yes you can give 3 medals to all 3 but firstly I must understand the solutions..

OpenStudy (anonymous):

\[\frac{\frac{1}{y} + \sin(\frac{1}{y})}{\frac{1}{y} + \cos(\frac{1}{y})}\]

hartnn (hartnn):

and y->0

OpenStudy (anonymous):

And limit y tends to 0 too..

hartnn (hartnn):

so u have now (1+ysin(1/y)) in numerator , right ?

OpenStudy (anonymous):

\[\frac{1 + y \sin(\frac{1}{y})}{1 + y \cos(\frac{1}{y})}\]

OpenStudy (anonymous):

There limit is going to 0 so 1/1 is 1.. Is this what you are saying @hartnn

OpenStudy (anonymous):

Or I said it in a hurry ??

hartnn (hartnn):

ok, now since sin fluctuates between 1 and -1 only, but y sin (...) will be very near to 0, because y is very near to 0 so, lim y->0 ysin(1/y) = 0

hartnn (hartnn):

similarly for denominator

OpenStudy (anonymous):

I liked it...but don't we consider sin(1/0)...its not defined, right?

OpenStudy (anonymous):

Yeah I got your approach too @hartnn Thanks for this help also...

OpenStudy (anonymous):

But we are multiplying it with a number nearer to 0...

hartnn (hartnn):

sin(1/0) is not defined but here y is NOT 0 its very very near to 0

OpenStudy (anonymous):

Excellent!!! got it.

hartnn (hartnn):

whoa! 8 medals given in this question!!

OpenStudy (anonymous):

Sorry not 8.. You have mistaken in a counting..

OpenStudy (anonymous):

u guys deserve more, very good explanations xD

hartnn (hartnn):

:D

OpenStudy (anonymous):

Total 9...

hartnn (hartnn):

yeah, 1 was given after i counted :P

OpenStudy (anonymous):

It is more like a Tutorial @hartnn

OpenStudy (anonymous):

Hahaha...yea 9..this is first time i've seen these many medals given!!!

hartnn (hartnn):

yo! 10 now. thnx cool. yes.

OpenStudy (anonymous):

lol yes

OpenStudy (anonymous):

tutorials should be like this, engaging way

OpenStudy (anonymous):

Superb discussion!!